OpenStudy (anonymous):

write how you would find the exact value of:

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ \sqrt{8} }\]

OpenStudy (austinl):

\(\dfrac{1}{2}\)

OpenStudy (anonymous):

write it out :)

OpenStudy (austinl):

How to get to the answer??

OpenStudy (anonymous):

ya

OpenStudy (austinl):

\(\dfrac{\sqrt{2}}{\sqrt{8}}\times \dfrac{\sqrt{8}}{\sqrt{8}}=\dfrac{\sqrt{16}}{8}=\dfrac{4}{8}=\dfrac{1}{2}\)

OpenStudy (anonymous):

thanks :-)

OpenStudy (austinl):

No problem.